A construction company can save $14,984 in Year 1 and increasing by $2,167 each year for the next 6 years by converting their diesel engines to biodiesel fuel. At in interest rate of 16% what is the equivalent annual worth of the savings?

*round to the nearest dollar without $ signs. For example, if your answer is $5678.09 enter 5678
Answer #1

Saving in first year = 14,984

This increases by 2,167 each year for the next 6 years

Life = 7 years

Interest Rate = 16%

What is the equivalent annual worth?

A = A1 + G (A/G, 16%, 7)

A = 14,984 + 2,167 (A/G, 16%, 7)

A = 14,984 + 2,167 (2.4169)

A = 20,221

This can be solved in alternative method.

Calculate PW

PW = A (P/A, 16%, 7) + G (P/G, 16%, 7)

PW = 14,984 (4.0386) + 2,167 (9.7610) = 81,666

Now calculate AW

AW = PW (A/P, 16%, 7)

AW = 81,666 (0.2476) = 20,221

The annual worth will be 20,221 (round to the nearest dollar)